OpenStudy (anonymous):

hat do you call a problem such as : solver the following for 'y' 3y-2/5 = 6y-5/11 i am trying to find a video to help me understand this

OpenStudy (mimi_x3):

Is it: \[\large \frac{3y-2}{5} = \frac{6y-5}{11} \]

OpenStudy (anonymous):

yes that is what it looks like

OpenStudy (mimi_x3):

Can you cross multiply it?

OpenStudy (anonymous):

i don't know how to figure it...

OpenStudy (mimi_x3):

\[(3y-2)*11\] for the first side. can you do the other side?

OpenStudy (anonymous):

(6y - 5) + 5 ?

OpenStudy (mimi_x3):

No, when you cross multipply its times. \[(6y-5)\times5\]

OpenStudy (anonymous):

so this may be stupid... but the = sign doesn't mean you do things separately?

OpenStudy (anonymous):

so basically you are just doing it as a cross multiply with the other problem, is that what you mean? the equal sign isn't going seperating the problems...

OpenStudy (mimi_x3):

Hm..its like this: \[11(3y-12)=5(6y-5)\] I was just trying to make it easier understand i failed i suppose xD

OpenStudy (anonymous):

oh no... it helps me to try to think it through... you were awesome!! ok, so i think i understand it... what exactly do you call this type of problem?

OpenStudy (mimi_x3):

Well, from that just expand it; and collect the like terms. Can you do that? Um..I forgot the name of it sorry.

OpenStudy (anonymous):

yes, i get it now! thank yoou soooo much!!!

OpenStudy (mimi_x3):

np. Just ask me if you need more help (: